OpenStudy (anonymous):

five less than one fourth the quantity of three times a number plus eight Let x = the number. A. B. C. D.

OpenStudy (anonymous):

\[\frac{1}{4} (3 x+8)-5=\frac{3 (x-4)}{4} \]